(4x^-2)-x^-1-5 put in quadratuc form to find x.
Answer by stanbon(75887)

(4x^-2)-x^-1-5 put in quadratic form to find x.
------
4(x^-1)^2 - (x^-1) - 5 = 0
---
To solve for "x" use the quadratic formula:
x^-1 = [1 +- sqrt(1-4*4*-5)]/8
---
x^-1 = [1 +- sqrt(81)]/8
x^-1 = [1 +- 9]/8
---
x^-1 = 10/8 or x = -1
---
So x = 4/5 or x = -1
============================
Cheers,
Stan H.
============================